# More Documents, Same Length: Isolating the Challenge of Multiple Documents in RAG

### v2

@@ 1 Introduction
Figure 1: More Documents, Same Length. We create various sets containing the same questions but differing in the number of distractor documents. Each set includes a multi-hop question, all of the supporting documents that contain the information to answer the question (pink), and varying distractor documents (blue).We begin with either 10 or 20 documents (depending on the dataset) as our full-doc version (left) and then reduce the number of documents while maintaining a fixed context size. When fewer documents are used, the remaining documents are extended (blue without text) so that concatenating them yields the same total length.
The RAG approach enriches prompts with relevant documents, retrieved according to an input query Karpukhin et al. (2020). For example, given a question about a certain historical period, RAG techniques can retrieve documents related to the time from a large historical corpus.
Recent work has noted a drop in RAG performance when retrieving many documents. For example, in multi-hop QA, LLMs struggle when the number of retrieved documents grows, even when presented with all the needed information Press et al. (2022); Liu et al. (2023); Levy et al. (2024); Wang et al. (2024). Such deficiencies were observed without controlling for the number of tokens in which the information is conveyed, i.e., when the number of documents grew, so did the number of overall tokens, thus conflating between the challenge of long context and multi document.
@@
In this work, we address the following question: Assuming a fixed input length, how is LLM performance affected by the number of retrieved documents? This disentangles the challenge of long context from the challenge in processing collections of related documents – which often contain redundancies, conflicting information, and implicit inter-document relations (Hirsch et al., 2023; Lior et al., 2024). From a practical perspective, answering this question can help understand a breadth versus depth tradeoff — i.e., whether to strive to retrieve shorter context out of many documents or whether to aim to retrieve longer context out of fewer documents. An ideal experimental setup would have the exact information conveyed in the same number of tokens across varying number of documents, from a long and self-contained single document to a large, multi-document corpus. We find that the custom sets we constructed from MuSiQue (Trivedi et al., 2022) and 2WikiMultiHopQA(Ho et al., 2020), a multi-hop QA datasets, serve as a convenient approximation, allowing us to explore the relationship between long-context and multi-document comprehension in a controlled environment with real-world texts.
We vary the number of documents in the input by gradually removing the distractor documents. When removing a distractor document, we respectively extend each of the remaining documents with distracting content from their corresponding Wikipedia article. Importantly, the process preserves the position of the relevant information within the cotext. This process is illustrated in Fig.˜1.
@@ 2 Multi-Document Evaluation with Controlled Token Count
Our datasets are based on MuSiQue (Trivedi et al., 2022) and 2WikiMultiHopQA (Ho et al., 2020), which we nickname as 2WMHQA. Both datasets are multi-hop QA datasets that consist of questions associated with paragraphs (20/10 paragraphs for MuSiQue/2WMHQA) sampled from individual documents, retrieved from Wikipedia according to the question. Of these paragraphs, 2–4 contain the supporting information necessary to answer the question, while the remaining paragraphs serve as realistic distractors in a RAG setup, as they are retrieved from related topics but do not contain relevant information to answer the question. Fig.˜1 shows an example query, and a list of retrieved documents, where three are relevant to the question (marked in pink), and the rest are distractors (marked in blue). We further elaborate on the dataset in section A in the appendix.
Leveraging MuSiQue’s and 2WMHQA’s structure, we constructed several data partitions to investigate the impact of the number of retrieved documents in a controlled manner. The process involved the following steps:
  1. 1.
    Select the total number of documents: We reduce the number of documents from the original document count down to only the supporting documents. For MuSiQue from 20 to 15, then 10, 8, and finally down to the 2–4 documents consisting of the relevant information to answer the question. Similarly for 2WMHQA, from 10 to 8, 4, and finally down to the 2 positive documents.
  2. 2.
    Choose the supporting and non-supporting documents: We always keep the documents that support the answer to ensure that the question remains answerable, and randomly select the remaining ones from the non-supporting set. Non-supporting documents remain consistent across different document counts, i.e., each set includes all documents from the smaller sets. Fig.˜1 shows such document selection in the two right columns, note that relevant documents (blue) are always kept.
  3. 3.
    Expand the selected documents: Since the original documents are Wikipedia paragraphs, we located their source Wikipedia pages and added text preceding and following the paragraphs to match the original token count. This replaces distracting content from different documents with distracting content from the same document. In Fig.˜1, we show that each of the remaining documents is expanded to keep the original token count, while ensuring that information from the supporting documents appeared in similar positions across all sets.
@@ 3.1 Experimental Setup
We evaluated six instruction-tuned LLMs from four model families: Llama-3.3 70B and Llama 3.2 3B  (AI@Meta, 2024)  [^2] A small counterpart to Llama-3.3 was not available at the time of evaluation., Qwen2.5 7B/72B (Qwen et al., 2025), Gemma2 9B/27B (Team et al., 2024), and GPT-4o/GPT-4o-mini (Hurst et al., 2024). Large models were run on Together.ai [^3] https://www.together.ai, and smaller ones on an A6000 GPU. We used a decoding temperature of 0.8, as recommended in prior evaluations (Chen et al., 2021). Evaluation relied on overlap F1 between gold and predicted outputs, following MuSiQue (Trivedi et al., 2022). Prompts, formats, and evaluation code were implemented using SEAM (Lior et al., 2024) (see Appendix C for details).
@@ 3.2 Results
Interestingly, GPT-4o performed significantly worse than GPT-4o-mini. Recent studies show GPT-4o-mini can outperform GPT-4o on certain tasks Chen et al. (2024); Nguyen et al. (2025); Alabbasi et al. (2025). This may be because GPT-4o’s larger parameter count leads to overfitting, while GPT-4o-mini’s smaller size forces it to focus on more generalizable patterns.
@@ Appendix A Datasets
We use two Multi-Hop QA datasets: MuSiQue (Trivedi et al., 2022) and 2WikiMultiHopQA (Ho et al., 2020), which we nickname as 2WMHQA. Both datasets consist of a set of questions associated with documents mined from Wikipedia. The documents are split between those that contain relevant knowledge to solve the question and distractors that contain similar details but not knowledge that is directly relevant to answering the question.
MuSiQue (Trivedi et al., 2022), a multi-hop QA dataset whose validation set consists of 2,417 answerable questions. Each question is associated with 20 paragraphs sampled from individual documents, retrieved from Wikipedia according to the question. Of these paragraphs, 2–4 contain the supporting information necessary to answer the question, while the remaining paragraphs serve as realistic distractors in a RAG setup, as they are retrieved from related topics but do not contain relevant information to answer the question. The mean token count for questions with their associated documents is 2,400 tokens per instance.
@@ B.3 Observed pattern applies to additional variants
We experimented with two additional model variants: Qwen-2 7B/72BYang et al. (2024) and Llama-3.1 8B/72BAI@Meta (2024). The observed trends remain consistent across different model versions. Qwen-2 demonstrates robustness as document count increases, suggesting it is better suited for multi-document processing, while Llama-3.1 shows a 10% performance decrease, as seen in Figure 4.
@@ B.4 Behavior is similar across instances with different amounts of tokens
We use prompt LABEL:lst:o1prompt for all model and document quantities taken from SEAM (Lior et al., 2024).

### v3

@@ 1 Introduction
Figure 1: We create various sets containing the same questions but differing in the number of distractor documents. Each set includes a multi-hop question, all of the supporting documents that contain the information to answer the question (pink), and varying distractor documents (blue). We begin with full-document version (left) and then reduce the number of documents while maintaining a fixed context size. When fewer documents are used, the remaining documents are extended so that concatenating them yields the same total length.
The RAG approach enriches prompts with relevant documents, retrieved according to an input query karpukhin-etal-2020-dense. For example, given a question about a certain historical period, RAG techniques can retrieve documents related to the time from a large historical corpus.
Recent work has noted a drop in RAG performance when retrieving many documents. For example, in multi-hop QA, LLMs struggle when the number of retrieved documents grows, even when presented with all the needed information press2022measuring; Liu2023LostIT; levy2024tasktokensimpactinput; wang2024leavedocumentbehindbenchmarking. Such deficiencies were observed without controlling for the number of tokens in which the information is conveyed, i.e., when the number of documents grew, so did the number of overall tokens, thus conflating between the challenge of long context and multi document.
In this work, we address the following question: Assuming a fixed input length, how is LLM performance affected by the number of retrieved documents? This disentangles the challenge of long context from the challenge in processing collections of related documents – which often contain redundancies, conflicting information, and implicit inter-document relations (hirsch-etal-2023-revisiting; lior2024seamstochasticbenchmarkmultidocument). From a practical perspective, answering this question can help understand a breadth versus depth tradeoff — i.e., whether to strive to retrieve shorter context out of many documents or whether to aim to retrieve longer context out of fewer documents. An ideal experimental setup would have the exact information conveyed in the same number of tokens across varying number of documents, from a long and self-contained single document to a large, multi-document corpus. We find that the custom sets we constructed from MuSiQue (trivedi2022musiquemultihopquestionssinglehop) and 2WikiMultiHopQA(ho2020constructing), a multi-hop QA datasets, serve as a convenient approximation, allowing us to explore the relationship between long-context and multi-document comprehension in a controlled environment with real-world texts.
As illustrated in Fig.˜1, we vary the number of documents in the input by gradually removing distractor documents. When removing a distractor document, we respectively extend each of the remaining documents with distracting content from their corresponding Wikipedia article. Importantly, the process preserves the position of the relevant information within the context.
@@ 2 Multi-Document Evaluation with Controlled Token Count
Our datasets are based on MuSiQue (trivedi2022musiquemultihopquestionssinglehop) and 2WikiMultiHopQA (ho2020constructing), which we nickname as 2WMHQA. Both datasets are multi-hop QA datasets that consist of questions associated with paragraphs (20/10 paragraphs for MuSiQue/2WMHQA) sampled from individual documents, retrieved from Wikipedia according to the question. Of these paragraphs, 2–4 contain the supporting information necessary to answer the question, while the remaining paragraphs serve as realistic distractors in a RAG setup, as they are retrieved from related topics but do not contain relevant information to answer the question. Fig.˜1 shows an example query, and a list of retrieved documents, where three are relevant to the question (marked in pink), and the rest are distractors (marked in blue). We further elaborate on the dataset in section A in the appendix.
Leveraging MuSiQue’s and 2WMHQA’s structure, we constructed several data partitions to investigate the impact of the number of retrieved documents in a controlled manner. The process involves the following three steps:
(1) Selecting the total number of documents:
We reduce the number of documents from the original document count down to only the supporting documents. For MuSiQue from 20 to 15, then 10, 8, and finally down to the 2–4 documents consisting of the relevant information to answer the question. Similarly for 2WMHQA, from 10 to 8, 4, and finally down to the 2 positive documents.
(2) Choosing the supporting and non-supporting documents:
We always keep the documents that support the answer to ensure that the question remains answerable, and randomly select the remaining ones from the non-supporting set. Non-supporting documents remain consistent across different document counts, i.e., each set includes all documents from the smaller sets. Fig.˜1 shows such document selection in the two right columns, note that relevant documents (blue) are always kept.
(3) Expanding the selected documents:
Since the original documents consist of Wikipedia paragraphs, we locate their source Wikipedia pages and add text preceding and following the paragraphs to match the original token count. Importantly, we uniformly expand supporting and non-supporting documents, under the assumption that both introduce distracting content, i.e., content which is not inherently needed to answer the question. In contrast, bianchi2025losthaystacksmallerneedles recently distinguishes between these two types of documents, and considers additional context from supporting documents as beneficial towards answering multi-hop questions. Future work may make finer-grained distinctions between document contexts by judging whether their contents help or distract with respect to individual questions.
@@ 3.1 Experimental Setup
We evaluated six instruction-tuned LLMs from four model families: Llama-3.3 70B and Llama 3.2 3B  (llama3modelcard)  [^2] A small counterpart to Llama-3.3 was not available at the time of evaluation., Qwen2.5 7B/72B (qwen2025qwen25technicalreport), Gemma2 9B/27B (team2024gemma), and GPT-4o/GPT-4o-mini (hurst2024gpt). Large models were run on Together.ai [^3] https://www.together.ai, and smaller ones on an A6000 GPU. We used a decoding temperature of 0.8, as recommended in prior evaluations (chen2021evaluating). Evaluation relied on overlap F1 between gold and predicted outputs, following MuSiQue (trivedi2022musiquemultihopquestionssinglehop). Prompts, formats, and evaluation code were implemented using SEAM (lior2024seamstochasticbenchmarkmultidocument) (see Appendix C for details).
@@ 3.2 Results
Interestingly, GPT-4o performed significantly worse than GPT-4o-mini. Recent studies show GPT-4o-mini can outperform GPT-4o on certain tasks chen2024steering; nguyen2025leveraging; alabbasi2025teleoracle. This may be because GPT-4o’s larger parameter count leads to overfitting, while GPT-4o-mini’s smaller size forces it to focus on more generalizable patterns.
@@ Appendix A Datasets
We use two Multi-Hop QA datasets: MuSiQue (trivedi2022musiquemultihopquestionssinglehop) and 2WikiMultiHopQA (ho2020constructing), which we nickname as 2WMHQA. Both datasets consist of a set of questions associated with documents mined from Wikipedia. The documents are split between those that contain relevant knowledge to solve the question and distractors that contain similar details but not knowledge that is directly relevant to answering the question.
MuSiQue (trivedi2022musiquemultihopquestionssinglehop), a multi-hop QA dataset whose validation set consists of 2,417 answerable questions. Each question is associated with 20 paragraphs sampled from individual documents, retrieved from Wikipedia according to the question. Of these paragraphs, 2–4 contain the supporting information necessary to answer the question, while the remaining paragraphs serve as realistic distractors in a RAG setup, as they are retrieved from related topics but do not contain relevant information to answer the question. The mean token count for questions with their associated documents is 2,400 tokens per instance.
@@ B.3 Observed pattern applies to additional variants
We experimented with two additional model variants: Qwen-2 7B/72Byang2024qwen2 and Llama-3.1 8B/72Bllama3modelcard. The observed trends remain consistent across different model versions. Qwen-2 demonstrates robustness as document count increases, suggesting it is better suited for multi-document processing, while Llama-3.1 shows a 10% performance decrease, as seen in Figure 4.
@@ B.4 Behavior is similar across instances with different amounts of tokens
We use prompt LABEL:lst:o1prompt for all model and document quantities taken from SEAM (lior2024seamstochasticbenchmarkmultidocument).